ANALYSIS OF THE HIGH-RESOLUTION MID-INFRARED SPECTRUM OF DEUTERATED WATER CLUSTERS

\underline{BRADLEY M. GIBSON} and JACOB T. STEWART, Department of Chemistry, University of Illinois at Urbana-Champaign, Urbana, IL 61801; BENJAMIN J. McCALL, Departments of Chemistry and Astronomy, University of Illinois at Urbana-Champaign, Urbana, IL 61801.

We have obtained rotationally-resolved infrared spectra of fully deuterated water clusters and Ar-D_{2}O clusters in the D_{2}O bending region near 1185-1205 cm^{−1} using a continuous-wave cavity ringdown spectrometer (cw-CRDS) with a quantum cascade laser (QCL) light source. These clusters were produced in a continuous supersonic expansion from a 150 μm x 1 cm slit using argon carrier gas, giving rovibrational transitions with a full width at half maximum of approximately 20 MHz. We will present our assignment of the Ar-D_{2}O spectra and preliminary analysis of spectra of other clusters (particularly our results for (D_{2}O)_{2}).